Part A Compute the energy separation between the ground and second excited states for an electron in a one-dimensional b...

Part A

Compute the energy separation between the ground and second excited states for an electron in a one-dimensional box that is 7.70 angstroms in length. Express the energy difference in kJ⋅mol−1.

Express your answer to three significant figures and include the appropriate units.

Part B

Compute the wavelength of light (in nm) corresponding to this energy.

Express your answer to three significant figures and include the appropriate units.
Answer #1

The energy of a particle in the nth level, according to the particle in a box model is given by:

En = 3.m. L

Where n is the level number, h is Planck's constant, m is the mass of the electron and L is the length of the box. The energy difference we are looking for is:

AB=B3-B1 = (91) 8. m. 12 8.14.,10-19 J (6.63.r10-34). ) m· L2 9.11x10-31 kg. (7.70x10-10m)2

In order to convert this to a wavelength, we need to use the relation:

h.c E = 1

Where c is the speed of light and lambda is the wavelength. We rearrange and solve:

hic = 6.63.c10-34 . s.3.x108m/s 10°m/s = 2.44c10-?M = 244nm 8.14., 10-19 E